OpenStudy (anonymous):

I am back once again. This math is not easy to me at all. Heeeellllppp 3y=4x+7 5y=4x+25 A. (2, 9) B. (5, 5) C. (5, 9) D. (2, 5)

OpenStudy (anonymous):

do simultaneously

OpenStudy (anonymous):

Hello Cupcake. we substract first from second via elimination method. 2y = 18 y = 18/2 = 9 5 * 9 = 4x + 25 45 = 4x + 25 4x = 20 x = 5. check: 3 * 9 = 4 * 5 + 7 27 = 27 Variant C(5 9)

OpenStudy (anonymous):

the ans is (c)